# RF-DETR: Neural Architecture Search for Real-Time Detection Transformers

### v1

@@ Abstract
Open-vocabulary detectors achieve impressive performance on COCO, but often fail to generalize to real-world datasets with out-of-distribution classes not typically found in their pre-training. Rather than simply fine-tuning a heavy-weight vision-language model (VLM) for new domains, we introduce RF-DETR, a light-weight specialist detection transformer that discovers accuracy-latency Pareto curves for any target dataset with weight-sharing neural architecture search (NAS). Our approach fine-tunes a pre-trained base network on a target dataset and evaluates thousands of network configurations with different accuracy-latency tradeoffs without re-training. Further, we revisit the “tunable knobs” for NAS to improve the transferability of DETRs to diverse target domains. Notably, RF-DETR significantly improves on prior state-of-the-art real-time methods on COCO and Roboflow100-VL. RF-DETR (nano) achieves 48.0 AP on COCO, beating D-FINE (nano) by 5.3 AP at similar latency, and RF-DETR (2x-large) outperforms GroundingDINO (tiny) by 1.2 AP on Roboflow100-VL while running $20\times$ as fast. To the best of our knowledge, RF-DETR (2x-large) is the first real-time detector to surpass 60 AP on COCO. Our code is available on GitHub.
@@ 1 Introduction
Object detection is a fundamental problem in computer vision that has matured in recent years (felzenszwalb2009object; lin2014coco; ren2015faster). Open-vocabulary detectors like GroundingDINO (liu2023grounding) and YOLO-World (Cheng2024YOLOWorld) achieve remarkable zero-shot performance on common categories like car, truck, and pedestrian. However, state-of-the-art vision-language models (VLMs) still struggle to generalize to out-of-distribution classes, tasks and imaging modalities not typically found in their pre-training (robicheaux2025roboflow100vl). Fine-tuning VLMs on a target dataset significantly improves in-domain performance at the cost of runtime efficiency (due to heavy-weight text encoders) and open-vocabulary generalization. In contrast, specialist (i.e., closed-vocabulary) object detectors like D-FINE (peng2024dfine) and RT-DETR (zhao2024rtdetr) achieve real-time inference, but underperform fined-tuned VLMs like GroundingDINO. In this paper, we modernize specialist detectors by combining internet-scale pre-training with real-time architectures to achieve state-of-the-art performance and fast inference.
@@
Are Specialist Detectors Over-Optimized for COCO? Sustained progress in object detection can be largely attributed to standardized benchmarks like PASCAL VOC (pascalvoc) and COCO (lin2014coco). However, we find that recent specialist detectors implicitly overfit to COCO at the cost of real-world performance using bespoke model architectures, learning rate schedulers, and augmentation schedulers. Notably, state-of-the-art object detectors like YOLOv8 (yolov8) generalize poorly to real-world datasets with significantly different data distributions from COCO (e.g., number of objects per image, number of classes, and dataset size). To address these limitations, we present RF-DETR, a scheduler-free approach that leverages internet-scale pre-training to generalize to real-world data distributions. To better specialize our model for diverse hardware platforms and dataset characteristics, we revisit neural architecture search (NAS) in the context of end-to-end object detection and segmentation.
@@
Rethinking Neural Architecture Search (NAS) for DETRs. NAS discovers accuracy-latency tradeoffs by exploring architectural variants within a pre-defined search space. NAS has been previously studied in the context of image classification (tan2019efficientnet; cai2019ofa) and for model sub-components like detector backbones tan2020efficientdet and FPNs ghiasi2019fpn. Unlike prior work, we explore end-to-end weight-sharing NAS for object detection and segmentation. Our key insight, inspired by OFA (cai2019ofa), is that we can vary model inputs like image resolution, and architectural components like patch size during training. Further, weight-sharing NAS allows us to modify inference configurations like the number of decoder layers and query tokens to specialize our strong base model without fine-tuning. We evaluate all model configurations with grid search on a validation set. Importantly, our approach does not evaluate the search space until the base model has been fully-trained on the target dataset. As a result, all possible sub-nets (i.e., model configurations within the search space) achieve strong performance without further fine-tuning, significantly reducing the computational cost of optimizing for new hardware. Interestingly, we find that sub-nets not explicitly seen during training still achieve high performance, suggesting that RF-DETR can generalize to unseen architectures (cf. H). Extending RF-DETR for segmentation is also relatively straightforward and only requires adding a lightweight instance segmentation head. We denote this model as RF-DETR-Seg. Notably, this allows us to also leverage end-to-end weight-sharing NAS to discover Pareto optimal architectures for real-time instance segmentation.
@@
Standardizing Latency Evaluation. We evaluate our approach on COCO (lin2014coco) and Roboflow100-VL (RF100-VL) (robicheaux2025roboflow100vl) and achieve state-of-the-art performance among real-time detectors. RF-DETR (nano) outperforms D-FINE (nano) by 5% AP on COCO at comparable run-times, and RF-DETR (2x-lage) beats GroundingDINO (tiny) on RF100-VL at a fraction of the runtime. RF-DETR-Seg (nano) outperforms YOLOv11-Seg (x-large) on COCO while running 4 $\times$ as fast. However, comparing RF-DETR’s latency with prior work remains challenging because reported latency evaluation varies significantly between papers. Notably, each new model re-benchmarks the latency of prior work for fair comparison on their hardware. For example, D-FINE’s reported latency evaluation of LW-DETR (chen2024lw) is 25% faster than originally reported. We identify that this lack of reproducibility can be primarily attributed to GPU power throttling during inference. We find that buffering between forward passes limits power over-draw and standardizes latency evaluation (cf. Table 1).
@@
Contributions. We present three major contributions. First, we introduce RF-DETR, a family of scheduler-free NAS-based detection and segmentation models that outperform prior state-of-the-art on RF100-VL (robicheaux2025roboflow100vl) and real-time methods with latencies $\leq$ 40 ms on COCO (lin2014coco)(cf. Fig. 1). To the best of our knowledge, RF-DETR is the first real-time detector to exceed 60 mAP on COCO. Next, we explore the “tunable-knobs” for weight-sharing NAS to improve accuracy-latency tradeoffs for end-to-end object detection (cf. Fig. 3). Notably, our use of a weight-sharing NAS allows us to leverage large-scale pre-training and effectively transfer to small datasets (cf. Tab. 4). Lastly, we revisit current benchmarking protocols for measuring latency and propose a simple standardized procedure to improve reproducibility.
@@ 2 Related Works
Neural Architecture Search (NAS) automatically identifies families of model architectures with different accuracy-latency tradeoffs (zoph2016neural; zoph2018learning; real2019regularized; cai2018efficient). Early NAS approaches (zoph2016neural; real2019regularized) focused primarily on maximizing accuracy, with little consideration for efficiency. As a result, discovered architectures (e.g., NASNet and AmoebaNet) were often computationally expensive. More recent hardware-aware NAS methods (cai2018proxylessnas; tan2019mnasnet; wu2019fbnet) address this limitation by incorporating hardware feedback directly into the search process. However, these methods must repeat the search and training process for each new hardware platform. In contrast, OFA (cai2019ofa) proposes a weight-sharing NAS that decouples training and search by simultaneously optimizing thousands of sub-nets with different accuracy-latency tradeoffs. Contemporary methods typically evaluate NAS for object detection by simply replacing standard backbones with NAS backbones in existing detection frameworks. Unlike prior work, we directly optimize end-to-end object detection accuracy to find Pareto optimal accuracy-latency tradeoffs for any target dataset.
@@
Real-Time Object Detectors are of significant interest for safety-critical and interactive applications. Historically, two-stage detectors like Mask-RCNN (he2017mask) and Hybrid Task Cascade (chen2019hybrid) achieved state-of-the-art performance at the cost of latency, while single-stage detectors like YOLO (redmon2016you) and SSD (liu2016ssd) traded accuracy for state-of-the-art runtime. However, modern detectors (zhao2024rtdetr) reexamine this accuracy-latency tradeoff, simultaneously improving on both axes. Recent YOLO variants innovate on architecture, data augmentation, and training techniques (redmon2016you; wang2023yolov7; wang2024yolov9; yolov8; yolov11) to improve performance while maintaining fast inference. Despite their efficiency, most YOLO models rely on non-maximum suppression (NMS), which introduces additional latency. In contrast, DETR (carion2020end) removes hand-crafted components like NMS and anchor boxes. However, early DETR variants (zhu2020deformable; zhang2022dino; meng2021conditional; liu2022dab) achieved strong accuracy at the cost of runtime, limiting their use in real-time applications. Recent works such as RT-DETR (zhao2024rtdetr) and LW-DETR (chen2024lw) have successfully adapted high performance DETRs for real-time applications.
@@
Vision-Language Models are trained on large-scale, weakly supervised image-text pairs from the web. Such internet-scale pre-training is a key enabler for open-vocabulary object detection (liu2023grounding; Cheng2024YOLOWorld). GLIP (li2021grounded) frames detection as phrase grounding with a single text query, while Detic (zhou2022detecting) boosts long-tail detection using ImageNet-level supervision (russakovsky2015imagenet). MQ-Det (xu2024multi) extends GLIP with a learnable module that enables multi-modal prompting. Recent VLMs demonstrate strong zero-shot performance and are often applied as “black-box” models in diverse downstream tasks (ma2023long; pmlr-v205-peri23a; khurana2024shelf; sal2024eccv; takmaz2025cal). However, robicheaux2025roboflow100vl find that such models perform poorly when evaluated on categories not typically found in their pre-training, requiring further fine-tuning. In addition, many vision-language models are prohibitively slow, making them difficult to use for real-time tasks. In contrast, RF-DETR combines the fast inference of real-time detectors with the internet-scale priors of VLMs to achieve state-of-the-art performance on RF100-VL and at all latencies $\leq$ 40 ms on COCO.
@@ 3 RF-DETR: Weight-Sharing NAS With Foundation Models
Figure 2: RF-DETR Architecture. RF-DETR uses a pre-trained ViT backbone to extract multi-scale features of the input image. We interleave windowed and non-windowed attention blocks to balance accuracy and latency. Notably, the deformable cross-attention layer and segmentation head both bilinearly interpolate the the output of the projector, allowing for consistent spatial organization of features. Lastly, we apply detection and segmentation losses at all decoder layers to facilitate decoder drop out at inference.
@@
In this section, we describe the architecture of our base model (cf. Fig. 2) and present the “tunable knobs” of our weight-sharing NAS (cf. Fig. 3). Further, we highlight the limitations of hand-designed learning-rate and augmentation schedulers, and advocate for a scheduler-free approach.
@@
Incorporating Internet-Scale Priors. RF-DETR modernizes LW-DETR (chen2024lw) by simplifying its architecture and training procedure to improve generalization to diverse target domains. First, we replace LW-DETR’s CAEv2 (zhang2022cae) backbone with DINOv2 (oquab2023dinov2). We find that initializing our backbone with DINOv2’s pre-trained weights significantly improves detection accuracy on small datasets. Notably, CAEv2’s encoder has 10 layers with a patch size of 16, while DINOv2’s encoder has 12 layers. Our DINOv2 backbone has more layers and is slower than CAEv2, but we make up for this latency using NAS (discussed next). Lastly, we facilitate training on consumer-grade GPUs via gradient accumulation by using layer norm instead of batch norm in the multi-scale projector.
@@
Real-Time Instance Segmentation. Inspired by li2023maskdino, we add a lightweight instance segmentation head to jointly predict high quality segmentation masks. Our segmentation head bilinearly interpolates the output of the encoder and learns a lightweight projector to generate a pixel embedding map. Specifically, we upsample the same low-resolution feature map for the detection and segmentation heads to ensure that it contains relevant spatial information. Unlike MaskDINO (li2023maskdino), we do not incorporate multi-scale backbone features in our segmentation head to minimize latency. Lastly, we compute the dot product of all projected query token embeddings (at the output of each decoder layer transformed by a FFN) with the pixel embedding map to generate segmentation masks. Interestingly, we can interpret these pixel embeddings as segmentation prototypes bolya2019yolact. Motivated by LW-DETR’s observation that pre-training improves DETRs, we pre-train RF-DETR-Seg on Objects-365 (objects365) psuedo-labeled with SAM2 (ravi2024sam2) instance masks.
@@
Figure 3: NAS Search Space. We vary (a) patch size, (b) number of decoder layers, (c) number of queries, (d) image resolution, and (e) number of windows per attention block in our weight-sharing NAS. In addition to training thousands of network configurations in parallel, we find that this “architecture augmentation” serves as a regularizer and improves generalization.
@@
End-to-End Neural Architecture Search. Our weight-sharing NAS evaluates thousands of model configurations with different input image resolutions, patch sizes, window attention blocks, decoder layers, and query tokens. At every training iteration, we uniformly sample a random model configuration and perform a gradient update. This allows our model to efficiently train thousands of sub-nets in parallel, similar to ensemble learning with dropout (dropout). We find that this weight-sharing NAS approach also serves as a regularizer during training, effectively performing “architecture augmentation”. To the best of our knowledge, RF-DETR is the first end-to-end weight-sharing NAS applied to object detection and segmentation. We describe each component below.
@@
  • Patch Size. Smaller patches lead to higher accuracy at greater computational cost. We adopt a FlexiVIT-style (beyer2023flexivit) transformation to interpolate between patch sizes during training.
@@
  • Number of Decoder Layers. Similar to recent DETRs (peng2024dfine; zhao2024rtdetr), we apply a regression loss to the output of all decoder layers during training. Therefore, we can drop any (or all) decoder blocks during inference. Interestingly, removing the entire decoder during inference effectively turns RF-DETR into a single-stage detector. Notably, truncating the decoder also shrinks the size of the segmentation branch, allowing for greater control over segmentation latency.
@@
At inference time, we pick a specific model configuration to select an operating point on the accuracy-latency Pareto curve. Importantly, different model configurations may have similar parameter counts but significantly different latencies. Similar to cai2019ofa, we see little benefit from fine-tuning the NAS-mined models on COCO (Appendix F), but note modest improvements from fine-tuning NAS-mined models on RF100-VL. We posit that RF-DETR on RF100-VL benefits from additional fine-tuning because the “architecture augmentation” regularization requires more than 100 epochs to converge on small datasets. Notably, prior weight-sharing NAS methods (cai2019ofa) train in stages and use a different learning-rate scheduler per-stage. However, such schedulers make strict assumptions about model convergence, which may not hold across diverse datasets.
Training Schedulers and Augmentations Bias Model Performance. State-of-the-art detectors often require careful hyper-parameter tuning to maximize performance on standard benchmarks. However, such bespoke training procedures implicitly bias the model towards certain dataset characteristics (e.g. number of images). Concurrent with DINOv3 (simeoni2025dinov3), we observe that cosine schedules assume a known (fixed) optimization horizon, which is impractical for diverse target datasets like those in RF100-VL. Data augmentations introduce similar biases by presuming prior knowledge of dataset properties. For example, prior work leverages aggressive data augmentation (e.g., VerticalFlip, RandomFlip, RandomResize, RandomCrop, YOLOXHSVRandomAug, and CachedMixUp) to increase effective dataset size. However, certain augmentations like VerticalFlip may negatively bias model predictions in safety-critical domains. For example, a person detector in a self-driving vehicle should not be trained with VerticalFlip to avoid false positive detections from reflections in puddles. Therefore, we limit augmentations to horizontal flips and random crops. Lastly, LW-DETR applies a per-image random resize augmentation, where each image is padded to match the largest image in the batch. As a result, most images have significant padding, which introduces window artifacts, and wastes computation on padded regions. In contrast, we resize images at the batch level to minimize the number of padded pixels per-batch and to ensure that all positional encoding resolutions are equally likely to be seen at train time.

### v2

@@ Abstract
Open-vocabulary detectors achieve impressive performance on COCO, but often fail to generalize to real-world datasets with out-of-distribution classes not typically found in their pre-training. Rather than simply fine-tuning a heavy-weight vision-language model (VLM) for new domains, we introduce RF-DETR, a light-weight specialist detection transformer that discovers accuracy-latency Pareto curves for any target dataset with weight-sharing neural architecture search (NAS). Our approach fine-tunes a pre-trained base network on a target dataset and evaluates thousands of network configurations with different accuracy-latency tradeoffs without re-training. Further, we revisit the “tunable knobs” for NAS to improve the transferability of DETRs to diverse target domains. Notably, RF-DETR significantly improves over prior state-of-the-art real-time methods on COCO and Roboflow100-VL. RF-DETR (nano) achieves 48.0 AP on COCO, beating D-FINE (nano) by 5.3 AP at similar latency, and RF-DETR (2x-large) outperforms GroundingDINO (tiny) by 1.2 AP on Roboflow100-VL while running $20\times$ as fast. To the best of our knowledge, RF-DETR (2x-large) is the first real-time detector to surpass 60 AP on COCO. Our code is available on GitHub.
@@ 1 Introduction
Object detection is a fundamental problem in computer vision that has matured in recent years (Felzenszwalb et al., 2009; Lin et al., 2014; Ren et al., 2015). Open-vocabulary detectors like GroundingDINO (Liu et al., 2023) and YOLO-World (Cheng et al., 2024) achieve remarkable zero-shot performance on common categories like car, truck, and pedestrian. However, state-of-the-art vision-language models (VLMs) still struggle to generalize to out-of-distribution classes, tasks and imaging modalities not typically found in their pre-training (Robicheaux et al., 2025). Fine-tuning VLMs on a target dataset significantly improves in-domain performance at the cost of runtime efficiency (due to heavy-weight text encoders) and open-vocabulary generalization. In contrast, specialist (i.e., closed-vocabulary) object detectors like D-FINE (Peng et al., 2024) and RT-DETR (Zhao et al., 2024) achieve real-time inference, but underperform fined-tuned VLMs like GroundingDINO. In this paper, we modernize specialist detectors by combining internet-scale pre-training with real-time architectures to achieve state-of-the-art performance and fast inference.
@@
Are Specialist Detectors Over-Optimized for COCO? Sustained progress in object detection can be largely attributed to standardized benchmarks like PASCAL VOC (Everingham et al., 2015) and COCO (Lin et al., 2014). However, we find that recent specialist detectors implicitly overfit to COCO at the cost of real-world performance using bespoke model architectures, learning rate schedulers, and augmentation schedulers. Notably, state-of-the-art object detectors like YOLOv8 (Jocher et al., 2023) generalize poorly to real-world datasets with significantly different data distributions from COCO (e.g., number of objects per image, number of classes, and dataset size). To address these limitations, we present RF-DETR, a scheduler-free approach that leverages internet-scale pre-training to generalize to real-world data distributions. To better specialize our model for diverse hardware platforms and dataset characteristics, we revisit neural architecture search (NAS) in the context of end-to-end object detection and segmentation.
@@
Rethinking Neural Architecture Search (NAS) for DETRs. NAS discovers accuracy-latency tradeoffs by exploring architectural variants within a pre-defined search space. NAS has been previously studied in the context of image classification (Tan and Le, 2019; Cai et al., 2019) and for model sub-components like detector backbones Tan et al. (2020) and FPNs Ghiasi et al. (2019). Unlike prior work, we explore end-to-end weight-sharing NAS for object detection and segmentation. Our key insight, inspired by OFA (Cai et al., 2019), is that we can vary model inputs like image resolution, and architectural components like patch size during training. Further, weight-sharing NAS allows us to modify inference configurations like the number of decoder layers and query tokens to specialize our strong base model without fine-tuning. We evaluate all model configurations with grid search on a validation set. Importantly, our approach does not evaluate the search space until the base model has been fully-trained on the target dataset. As a result, all possible sub-nets (i.e., model configurations within the search space) achieve strong performance without further fine-tuning, significantly reducing the computational cost of optimizing for new hardware. Interestingly, we find that sub-nets not explicitly seen during training still achieve high performance (Appendix F), suggesting that RF-DETR can generalize to unseen architectures. Extending RF-DETR for segmentation is also relatively straightforward and only requires adding a lightweight instance segmentation head. We denote this model as RF-DETR-Seg. Notably, this allows us to also leverage end-to-end weight-sharing NAS to discover Pareto optimal architectures for real-time instance segmentation.
@@
Standardizing Latency Evaluation. We evaluate our approach on COCO (Lin et al., 2014) and Roboflow100-VL (RF100-VL) (Robicheaux et al., 2025) and achieve state-of-the-art performance among real-time detectors. RF-DETR (nano) outperforms D-FINE (nano) by 5% AP on COCO at comparable run-times, and RF-DETR (2x-large) beats GroundingDINO (tiny) on RF100-VL at a fraction of the runtime. RF-DETR-Seg (nano) outperforms YOLOv11-Seg (x-large) on COCO while running 4 $\times$ as fast. However, comparing RF-DETR’s latency with prior work remains challenging because reported latency evaluation varies significantly between papers. Notably, each new model re-benchmarks the latency of prior work for fair comparison on their hardware. For example, D-FINE’s reported latency evaluation of LW-DETR (Chen et al., 2024a) is 25% faster than originally reported. We identify that this lack of reproducibility can be primarily attributed to GPU power throttling during inference. We find that buffering between forward passes limits power over-draw and standardizes latency evaluation (Table 1).
@@
Contributions. We present three major contributions. First, we introduce RF-DETR, a family of scheduler-free NAS-based detection and segmentation models that outperform prior state-of-the-art on RF100-VL (Robicheaux et al., 2025) and real-time methods with latencies $\leq$ 40 ms on COCO (Lin et al., 2014) (Figure 1). To the best of our knowledge, RF-DETR is the first real-time detector to exceed 60 mAP on COCO. Next, we explore the “tunable-knobs” for weight-sharing NAS to improve accuracy-latency tradeoffs for end-to-end object detection (Figure 3). Notably, our use of a weight-sharing NAS allows us to leverage large-scale pre-training and effectively transfer to small datasets (Table 4). Lastly, we revisit current benchmarking protocols for measuring latency and propose a simple standardized procedure to improve reproducibility.
@@ 2 Related Works
Neural Architecture Search (NAS) automatically identifies families of model architectures with different accuracy-latency tradeoffs (Zoph and Le, 2016; Zoph et al., 2018; Real et al., 2019; Cai et al., 2018a). Early NAS approaches (Zoph and Le, 2016; Real et al., 2019) focused primarily on maximizing accuracy, with little consideration for efficiency. As a result, discovered architectures (e.g., NASNet and AmoebaNet) were often computationally expensive. More recent hardware-aware NAS methods (Cai et al., 2018b; Tan et al., 2019; Wu et al., 2019) address this limitation by incorporating hardware feedback directly into the search process. However, these methods must repeat the search and training process for each new hardware platform. In contrast, OFA (Cai et al., 2019) proposes a weight-sharing NAS that decouples training and search by simultaneously optimizing thousands of sub-nets with different accuracy-latency tradeoffs. Contemporary methods typically evaluate NAS for object detection by simply replacing standard backbones with NAS backbones in existing detection frameworks. Unlike prior work, we directly optimize end-to-end object detection accuracy to find Pareto optimal accuracy-latency tradeoffs for any target dataset.
@@
Real-Time Object Detectors are of significant interest for safety-critical and interactive applications. Historically, two-stage detectors like Mask-RCNN (He et al., 2017) and Hybrid Task Cascade (Chen et al., 2019) achieved state-of-the-art performance at the cost of latency, while single-stage detectors like YOLO (Redmon et al., 2016) and SSD (Liu et al., 2016) traded accuracy for state-of-the-art runtime. However, modern detectors (Zhao et al., 2024) reexamine this accuracy-latency tradeoff, simultaneously improving on both axes. Recent YOLO variants innovate on architecture, data augmentation, and training techniques (Redmon et al., 2016; Wang et al., 2023; 2024; Jocher et al., 2023; 2024) to improve performance while maintaining fast inference. Despite their efficiency, most YOLO models rely on non-maximum suppression (NMS), which introduces additional latency. In contrast, DETR (Carion et al., 2020) removes hand-crafted components like NMS and anchor boxes. However, early DETR variants (Zhu et al., 2020; Zhang et al., 2022a; Meng et al., 2021; Liu et al., 2022) achieved strong accuracy at the cost of runtime, limiting their use in real-time applications. Recent works such as RT-DETR (Zhao et al., 2024) and LW-DETR (Chen et al., 2024a) have successfully adapted high performance DETRs for real-time applications. Building on LW-DETR, RF-DETR is the first real-time detector to achieve more than 60 AP on COCO.
@@
Vision-Language Models are trained on large-scale, weakly supervised image-text pairs from the web. Such internet-scale pre-training is a key enabler for open-vocabulary object detection (Liu et al., 2023; Cheng et al., 2024). GLIP (Li et al., 2022) frames detection as phrase grounding with a single text query, while Detic (Zhou et al., 2022) boosts long-tail detection using ImageNet-level supervision (Russakovsky et al., 2015). MQ-Det (Xu et al., 2024) extends GLIP with a learnable module that enables multi-modal prompting. Recent VLMs demonstrate strong zero-shot performance and are often applied as “black-box” models in diverse downstream tasks (Ma et al., 2023; Peri et al., 2023; Khurana et al., 2024; Osep et al., 2024; Takmaz et al., 2025). However, Robicheaux et al. (2025) find that such models perform poorly when evaluated on categories not typically found in their pre-training, requiring further fine-tuning. In addition, many vision-language models are prohibitively slow, making them difficult to use for real-time tasks. In contrast, RF-DETR combines the fast inference of real-time detectors with the internet-scale priors of VLMs to achieve state-of-the-art performance on RF100-VL and at all latencies $\leq$ 40 ms on COCO.
@@ 3 RF-DETR: Weight-Sharing NAS With Foundation Models
Figure 2: RF-DETR Architecture. RF-DETR uses a pre-trained ViT backbone to extract multi-scale features of the input image. We interleave windowed and non-windowed attention blocks to balance accuracy and latency. Notably, the deformable cross-attention layer and segmentation head both bilinearly interpolate the output of the projector, allowing for consistent spatial organization of features. Lastly, we apply detection and segmentation losses at all decoder layers to facilitate decoder drop out at inference.
@@
In this section, we describe the architecture of our base model (Figure 2) and present the “tunable knobs” of our weight-sharing NAS (Figure 3). Further, we highlight the limitations of hand-designed learning-rate and augmentation schedulers, and advocate for a scheduler-free approach.
@@
Incorporating Internet-Scale Priors. RF-DETR modernizes LW-DETR (Chen et al., 2024a) by simplifying its architecture and training procedure to improve generalization to diverse target domains. First, we replace LW-DETR’s CAEv2 (Zhang et al., 2022b) backbone with DINOv2 (Oquab et al., 2023). We find that initializing our backbone with DINOv2’s pre-trained weights significantly improves detection accuracy on small datasets. Notably, CAEv2’s encoder has 10 layers with a patch size of 16, while DINOv2’s encoder has 12 layers. Our DINOv2 backbone has more layers and is slower than CAEv2, but we make up for this latency using NAS (discussed next). Lastly, we facilitate training on consumer-grade GPUs via gradient accumulation by using layer norm instead of batch norm in the multi-scale projector.
@@
Real-Time Instance Segmentation. Inspired by Li et al. (2023), we add a lightweight instance segmentation head to jointly predict high quality segmentation masks. Our segmentation head bilinearly interpolates the output of the encoder and learns a lightweight projector to generate a pixel embedding map. Specifically, we upsample the same low-resolution feature map for the detection and segmentation heads to ensure that it contains relevant spatial information. Unlike MaskDINO (Li et al., 2023), we do not incorporate multi-scale backbone features in our segmentation head to minimize latency. Lastly, we compute the dot product of all projected query token embeddings (at the output of each decoder layer transformed by a FFN) with the pixel embedding map to generate segmentation masks. Interestingly, we can interpret these pixel embeddings as segmentation prototypes (Bolya et al., 2019). Motivated by LW-DETR’s observation that pre-training improves DETRs, we pre-train RF-DETR-Seg on Objects-365 (Shao et al., 2019) psuedo-labeled with SAM2 (Ravi et al., 2024) instance masks.
@@
Figure 3: NAS Search Space. We vary (a) patch size, (b) number of decoder layers, (c) number of queries, (d) image resolution, and (e) number of windows per attention block when evaluating different operating points along RF-DETR’s Pareto curve. In addition to training thousands of network configurations in parallel, we find that this “architecture augmentation” serves as a regularizer and improves generalization.
@@
End-to-End Neural Architecture Search. Our weight-sharing NAS evaluates thousands of model configurations with different input image resolutions, patch sizes, window attention blocks, decoder layers, and query tokens. At every training iteration, we uniformly sample a random model configuration and perform a gradient update (Appendix A). This allows our model to efficiently train thousands of sub-nets in parallel, similar to ensemble learning with dropout (Srivastava et al., 2014). We find that this weight-sharing NAS approach also serves as a regularizer during training, effectively performing “architecture augmentation”. To the best of our knowledge, RF-DETR is the first end-to-end weight-sharing NAS applied to object detection and segmentation. We describe each component below.
@@
  • Patch Size. Smaller patches lead to higher accuracy at greater computational cost. We adopt a FlexiVIT-style (Beyer et al., 2023) transformation to interpolate between patch sizes during training.
@@
  • Number of Decoder Layers. Similar to recent DETRs (Peng et al., 2024; Zhao et al., 2024), we apply a regression loss to the output of all decoder layers during training. Therefore, we can drop any (or all) decoder blocks during inference. Interestingly, removing the entire decoder during inference effectively turns RF-DETR into a single-stage detector. Notably, truncating the decoder also shrinks the size of the segmentation branch, allowing for greater control over segmentation latency.
@@
At inference time, we pick a specific model configuration to select an operating point on the accuracy-latency Pareto curve. Importantly, different model configurations may have similar parameter counts but significantly different latencies. Similar to Cai et al. (2019), we see little benefit from fine-tuning the NAS-mined models on COCO (Appendix G), but note modest improvements from fine-tuning NAS-mined models on RF100-VL. This additional fine-tuning is optional, and is often unnecessary for practical deployment. We posit that RF-DETR benefits from additional fine-tuning on RF100-VL because the “architecture augmentation” regularization requires more than 100 epochs to converge on small datasets. Notably, prior weight-sharing NAS methods (Cai et al., 2019) train in stages and use a different learning-rate scheduler per-stage. However, such schedulers make strict assumptions about model convergence, which may not hold across diverse datasets.
Training Schedulers and Augmentations Bias Model Performance. State-of-the-art detectors often require careful hyper-parameter tuning to maximize performance on standard benchmarks. However, such bespoke training procedures implicitly bias the model towards certain dataset characteristics (e.g. number of images). Concurrent with DINOv3 (Siméoni et al., 2025), we observe that cosine schedules assume a known (fixed) optimization horizon, which is impractical for diverse target datasets like those in RF100-VL. Data augmentations introduce similar biases by presuming prior knowledge of dataset properties. For example, prior work leverages aggressive data augmentation (e.g., VerticalFlip, RandomFlip, RandomResize, RandomCrop, YOLOXHSVRandomAug, and CachedMixUp) to increase effective dataset size. However, certain augmentations like VerticalFlip may negatively bias model predictions in safety-critical domains. For example, a person detector in a self-driving vehicle should not be trained with VerticalFlip to avoid false positive detections from reflections in puddles. Therefore, we limit augmentations to horizontal flips and random crops. Lastly, LW-DETR applies a per-image random resize augmentation, where each image is padded to match the largest image in the batch. As a result, most images have significant padding, which introduces window artifacts, and wastes computation on padded regions. In contrast, we resize images at the batch level to minimize the number of padded pixels per-batch and to ensure that all positional encoding resolutions are equally likely to be seen at train time.